OpenStudy (anonymous):

Give an example of a rational function that has no horizontal asymptote and a vertical asymptote at x = 1.

Parth (parthkohli):

That means that \(x = 1\) is an excluded value.

OpenStudy (anonymous):

im not really sure what u mean

Parth (parthkohli):

That means that we won't have \(x = 1 \) in the domain.\[{x^2 \over x - 2} \]Has a vertical asymptote at \(x = 2\).

OpenStudy (anonymous):

thats it?

Parth (parthkohli):

Yep, just that you have to find it for \(x = 1\).

OpenStudy (anonymous):

so i can just put x^2/x-2

Parth (parthkohli):

Nope.

Parth (parthkohli):

Hint: You can't input 1 in\[x^2\over x - 1 \]

OpenStudy (anonymous):

im lost at this point

OpenStudy (anonymous):

Parth was giving you an example of the concept, and then hoping that you would then apply it to your particular problem.

OpenStudy (anonymous):

right but I'm not really understanding how to do the question i need help finding the example of what they are asking for

Parth (parthkohli):

Do you know what the 'excluded value' means?

OpenStudy (anonymous):

Excluded values are simply that: values that are excluded, or left out. These are values that will make the denominator of a rational expression equal to 0.

Parth (parthkohli):

Exactly. The question merely means that the excluded value of \(x\) is \(1\).

OpenStudy (anonymous):

Well there you go. Make an equation such that the denominator is equal to 0 when x=1, and then just make sure that it doesn't have a horizontal asymptote.

OpenStudy (asnaseer):

you could look at it in this way:\[f(x)=\frac{g(x)}{h(x)}\]find a function \(h(x)\) such that \(h(1)=0\)

OpenStudy (anonymous):

4x^4y^3 ______ -8x^2y

Parth (parthkohli):

Woo! That's a great explanation, but it's a little complicating for a beginner.

Parth (parthkohli):

KISS principle: http://en.wikipedia.org/wiki/KISS_principle

OpenStudy (asnaseer):

keep it simple :)

Parth (parthkohli):

You can even make it something like \[x\over x - 3 \]If 3 is the horizontal asymptote :P

Parth (parthkohli):

@asnaseer That KISS Principle link was for @babydoll332 haha She posted a little too complicated expression.

OpenStudy (asnaseer):

yes I /guessed/ it was :)

OpenStudy (asnaseer):

@babydoll332 maybe a different approach might help. if:\[y=x-a\]where a is some constant, then what value of x will make y zero?

Parth (parthkohli):

Just wanted to make sure that you didn't take that. All swell! :D

OpenStudy (anonymous):

so is my expression right or wrong or is it to complex for u to explain that @ParthKohli

OpenStudy (asnaseer):

your expression is wrong

Parth (parthkohli):

The point is that there are flaws in your expression - and also you need to keep your answers simple! :)

OpenStudy (asnaseer):

and far too complex than required

Parth (parthkohli):

@asnaseer Congratulations for becoming a champ!

OpenStudy (asnaseer):

?

OpenStudy (anonymous):

ok then give me an easier example so i can make it the way i want after

Parth (parthkohli):

You're a champion at problem solving haha

Parth (parthkohli):

I did give one: If x = 3 is a vertical asymptote, then we can make the function look like\[f(x) = {x \over x - 3} \]

OpenStudy (anonymous):

thats the it thats the whole thing way the question is asking for

Parth (parthkohli):

Because then, plugging 3 into \(x\) will make the denominator 0. You did post how the denominator is 0 in a vertical asymptote, didn't you?

Parth (parthkohli):

You have to make an expression where plugging 1 in \(x\) makes the denominator 0. All right?

OpenStudy (asnaseer):

Parth - you also need to ensure it has no horizontal symptotes

OpenStudy (anonymous):

u are confusing me omg !!!

OpenStudy (asnaseer):

@babydoll332 sorry if my comments confused you further. maybe I will let Parth explain this to you.

Parth (parthkohli):

I was going to leave it to you >.>

OpenStudy (anonymous):

nooo i want @asnaseer

Parth (parthkohli):

Maybe it's my words that confused her.

OpenStudy (asnaseer):

I don't mind trying to explain it as long as my friend Parth agrees. :)

Parth (parthkohli):

Continue, I am watching you guys\[\Huge \ddot \smile \]

OpenStudy (anonymous):

this is my example x+1 /x(x+4)

Parth (parthkohli):

Keep in mind: Does plugging in 1 make the denominator 0?

OpenStudy (asnaseer):

ok, now if we take your example, we see that at x=1 we get:\[\frac{1+1}{1(1+4)}=\frac{2}{5}\ne\infty\]

OpenStudy (asnaseer):

you need to find an equation such that when you set x=1, the denominator will become zero.

OpenStudy (asnaseer):

so lets first look at a simple function. y = x - a what value of x makes y zero?

OpenStudy (anonymous):

(x-1)(x+2)

OpenStudy (asnaseer):

yes that would work as the denominator

OpenStudy (asnaseer):

so now you know that you have something of the form:\[y=\frac{g(x)}{(x-1)(x+2)}\]where g(x) is some other function of x. the only problem with the example you have selected is that it makes the denominator equal to zero at x=1 AND at x=-2

OpenStudy (asnaseer):

so this has TWO vertical asymptotes - which is not what was asked for. Can you think of a simpler example?

OpenStudy (anonymous):

(x-1)(x-2)?

OpenStudy (asnaseer):

that still has TWO vertical asymptotes - one at x=1 and the other at x=2

OpenStudy (asnaseer):

think even simpler

OpenStudy (anonymous):

so (x+1)(x+2)?

OpenStudy (asnaseer):

that will have asymptotes at x=-1 and x=-2 but you were asked to find one that has an asymptote at x=1 it looks like you are trying to get some quadratic equation to fit this question. remember it does not have to be quadratic. think of a polynomial of degree 1.

OpenStudy (asnaseer):

does that make sense?

OpenStudy (asnaseer):

a polynomial of degree 1 would be of the form:\[ax+b\]where a and b are constants.

OpenStudy (anonymous):

(x+1)/(x^2-4)

Parth (parthkohli):

Remember that x + 1 should be the denominator.

Parth (parthkohli):

You must get the denominator as 0, right?

OpenStudy (anonymous):

i guess i don't know thats why i am asking because i don't get the question nor what the example would be

OpenStudy (asnaseer):

the example you just gave doesn't work as it makes the denominator zero when x=2 or x=-2

OpenStudy (anonymous):

Reread the explanations that asnaseer and parth have both given you. Then reread the section of your textbook that talks about this subject. Then think about the subject for ten minutes. Then come back, and if you still don't get it, ask specific questions about specific parts that you do not understand.

OpenStudy (asnaseer):

let me try one more approach first..

OpenStudy (anonymous):

i just want an example thats all I'm asking for so then i can go and actually apply it and understand what both of you are trying to explain

OpenStudy (asnaseer):

1) do you agree that you will get a vertical asymptote if you have a rational function where the denominator can become zero?

OpenStudy (anonymous):

They've given you roughly a dozen examples by now.

OpenStudy (anonymous):

ok yea

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):

good, now, in order to get a vertical asymptote at x=1, you need to find a function in the denominator that will become zero when x=1. agreed?

OpenStudy (anonymous):

agreeed

OpenStudy (asnaseer):

great - we are ALMOST there now...

OpenStudy (asnaseer):

what is the SIMPLEST function you can think of that will become zero when x=1?

OpenStudy (anonymous):

x=o

OpenStudy (asnaseer):

that is just stating x=0 you need to create a function that, when you plug x=1 into it, will give you a value of zero.

OpenStudy (asnaseer):

e.g. if the function were y = x - 3 then this gives a value of zero when x=3

OpenStudy (asnaseer):

because at x=3 you get y = 3 - 3 = 0

OpenStudy (anonymous):

ok so y=x-1

OpenStudy (asnaseer):

perfect!

OpenStudy (asnaseer):

so we have now defined the function for the denominator - it is (x-1)

OpenStudy (asnaseer):

the next step is to find a function for the numerator such that we do NOT get any horizontal asymptotes

OpenStudy (asnaseer):

do you know what properties of a function will give a horizontal asymptote?

OpenStudy (anonymous):

not really

OpenStudy (asnaseer):

ok, let me try to explain it...

OpenStudy (asnaseer):

we know your final function has to be of the form:\[y=\frac{g(x)}{x-1}\]where g(x) is the function we need to find for the numerator. now, if y tends to some fixed value as x tends to plus/minus infinity, then this means that y has a horizontal asymptote

OpenStudy (anonymous):

I think the easier explanation for someone at this level might have to use graphing =/

OpenStudy (asnaseer):

e.g. lets say you had:\[y=\frac{x}{x-1}\]then can you see that as x tends to infinity then y tends to 1?

OpenStudy (asnaseer):

good point nbouscal - let me draw a graph...

OpenStudy (anonymous):

I'm good i don't need to see a graph

OpenStudy (anonymous):

How are you going to solve it, then?

OpenStudy (asnaseer):

are you sure babydoll332?

OpenStudy (anonymous):

ya i think I'm good

OpenStudy (anonymous):

Ok, what is the solution then??